Identify the level curves of the following functions:

(i) √(x2+y2)

(ii) √(4 - xy2)

(iii) x-y

(iv) x/y



1
Expert's answer
2022-01-10T17:43:28-0500

i)

"\sqrt{x^2+y^2}=k"

"x^2+y^2=k^2"

circles with centre at origin and radius k


ii)

"\sqrt{4-x^2+y^2}=k"

"\frac{y^2}{k^2-4}-\frac{x^2}{k^2-4}=1"

hyperbola with asymptotes "y=\pm x"


iii)

"x-y=k"

line with slope 1


iv)

"x\/y=k"

"y=x\/k"

line with slope "1\/k"